Correlator polB sbc 1 gain drifts

march 04

a1705 up to 15apr03
list of all files with drifts 15apr03 to 27may03

Resolution: 01jun03

    The 260->baseband mixer chassis was tested on the bench and the tubular lowpass filters (k&E)  (100 mhz) after the mixer were causing the problem. We cooled all 8 tubular filters and 1b,4a,4b showed large gain variations. Removing these filters, we saw that the impedance between the center conductor and ground varied from 1 Mohm to 2 ohms. Opening the filters we found that the mylar sheet that was wrapped around the filter core and acted as the dielectric material was causing the problem. It is a flat sheet that wraps around the filter core leaving a thin line where the filter could be tuned.  At the ends of the filter, the gap got too large and the filter core (actually the capacitor lugs) were getting close the outer case and shorting. As the temperature changed, the impedence would change and cause the gain variations that we were seeing. It looks like this is a basic design flaw with the filters. We ordered a new set of 8 filters and have now become suspicious of these "inactive" devices. All 8 filters were replaced on 27jun03.

a1705 and if2 noise source (thur 15apr03)

 Experiment A1705 was experiencing drifts in polb sbc 1 of their data. It was centered near 1400 Mhz.
The plots show the 0 lag power for the entire night. The large jumps are when the attenuators were changed. The narrow spikes are the cals. The slow variation that occurs in 1 pol and not the other is the problem. The gradual rise in all sbc is continuum.

A1705 data

    This is the data from A1705.

Noise source data:

    This data was taken with the if2 noise source inserted. It is 1 second dumps just like the a1705 data. Here is some of the evidence that has been gleaned(in time order)
processing: usr/a1705/

After 15apr03

    The problem went away and then reappeared. Here is a list of all of the observations that have had drifts 15apr03 thru 27may03. There are 123 files that were checked. 23 files had gain drifts. Some of the files measured one polarization per board so we would never see a  1b drift in their data (eg. a1708).
file drift
corfile.09apr03.if2noise.1 1b
corfile.10apr03.if2noise.1 1b (1)
corfile.14apr03.if2noise.1 1b (2)
corfile.09may03.x101.3 4a
corfile.12may03.a1312.1 1b
corfile.13may03.a1457.1 1b
corfile.15may03.a1708.1 4a
corfile.16may03.a1703.2 4a
corfile.16may03.a1708.1 4a
corfile.16may03.reu03.2 1b
corfile.17may03.a1312.2 4a
corfile.18may03.a1703.1 4a (3)
corfile.19may03.a1708.2 4a
corfile.19may03.if2noise.1 4a
corfile.20may03.a1708.1 4a
corfile.23may03.a1703.1 4a
corfile.24may03.a1545.1 1b
corfile.24may03.a1703.1 4a,1b(3)
corfile.24may03.a1789.1 4a
corfile.24may03.x111.1 4a
corfile.25may03.a1703.1 1b,4a(3)
corfile.25may03.a1708.1 4a
corfile.26may03.a1703.1 1b*
  1. if2 rack outputs  1b,2b switched. if2_1b -> corInp_2b, if2_2b->corinp_1b. The switch was done at the output of if2 rack. The problem stayed in cor_1b, so the problem is not in the if2 racks.
  2. The correlator attenenuator chassis outputs attn_1b,attn_2b were switched at the output of the atttenuator chassis. The problem stayed in _1b so the problem is downstream from the attenuator box (cable, mixer chassis, etc). The power counters also drift so it is before the split of the power counters in the mixer chassis..
  3. A large fraction of the datafile had gain drifts...